Solve by substitution
x = y + 5
2x - 2y = 12
Please show me if this problem is independent, dependent or inconsistent. More importantly I'd like to know the best way to solve future problems like this''.

Answer by ewatrrr(24785)


Hi,
Solving by substitution
x = y + 5
2x - 2y = 12 |substituting (y+5) for x
2(y+5) - 2y = 12
2y + 10 - 2y = 12
10 ≠ 12 there is NO Solution
This is an Inconsistent System of Equations